Find g 0 f.
f(x)=x+5
g(x)=x^2

Answer by stanbon(75887)

Find g 0 f.
f(x)=x+5
g(x)=x^2
-----
gof
= g[f(x)]
= g[x+5]
= (x+5)^2
= x^2+10x+25
=================================================
Cheers,
Stan H.
===============